The final mathematics exam at a certain school contains question of equal difficulty, and is matched to the students' past performance in mathematics. Four students who take the exam have averages of 25%, 50%, 60% and 80%. What is the probability that at least one of these students will get the first question on the exam correct?
A)97/100 B)9/10 C)17/20 D)13/15 E)23/30

Answer by greenestamps(13330)



P(at least one gets the first question correct) = 1 - P(none get it correct)

P(1 wrong) = 75% = 3/4
P(2 wrong) = 50% = 1/2
P(3 wrong) = 40% = 2/5
P(4 wrong) = 20% = 1/5

P(none get it correct) = (3/4)(1/2)(2/5)(1/5) = 6/200 = 3/100

P(at least one gets it correct = 1-3/100 = 97/100

ANSWER A